i know it said don't dump your homework but i only have these two for college algebra. please get back
thank you
An executive invests $22,000, some at 7% and some at 5% annual interest. If he receives an annual return of $1,420, how much is invested at each rate?
A doctor prescribes an ointment that is 4% hydrocortisone. A pharmacist has 1% and 5% concentrations in stock. How many ounces of each should the pharmacist use to make a 1-ounce tube?

Answer by mananth(16946)

Part I 7.00% per annum ------------- Amount invested =x
Part II 5.00% per annum ------------ Amount invested = y
22000
Interest----- 1420.00

Part I 7.00% per annum ---x
Part II 5.00% per annum ---y
Total investment
x + 1 y= 22000 -------------1
Interest on both investments
7.00% x + 5.00% y= 1420
Multiply by 100
7 x + 5 y= 142000.00 --------2
Multiply (1) by -7
we get
-7 x -7 y= -154000.00
Add this to (2)
0 x -2 y= -12000
divide by -2
y = 6000
Part I 7.00% $ 16000
Part II 5.00% $ 6000

CHECK
16000 --------- 7.00% ------- 1120.00
6000 ------------- 5.00% ------- 300.00
Total -------------------- 1420.00
------------------
Hydrocortisone
percent ---------------- quantity
Type I 1.00% ---------------- x ounce
Type II 5.00% ------ 1 - x ounce
Mixture 4.00% ---------------- 1
Total 1 ounce
1.00% x + 5.00% ( 1 - x ) = 4.00% * 1
1 x + 5 ( 1 - x ) = 4
1 x + 5 - 5 x = 4
1 x - 5 x = 4 - -5
-4 x = -1
/ -4
x = 0.25 ounce 1.00% Type I
0.75 ounce 5.00% Type II



m.ananth@hotmail.ca